Solve Problems Using Linear Systems
8. Hakim invested $15 000. He put part of it in a term deposit that paid 4 percent per annum, and the remainder in a treasury bill that paid 5 percent per annum. After one year, the total interest was $690. How much did Hakim invest each rate?
can you please help me ..i dont understand it thank u

Answer by Earlsdon(6294)

Let x = the amount that Hakim invested at 4% per annum, then the rest ($15,000-x) was invested at 5% per annum.
The interest earned on the Term Deposit at 4% per annum can be expressed as 0.04x while the interest earned on the Treasury Bill at 5% can be expressed as 0.05($15,000-x) and the sum of these two amounts is given as $690.
You can now form the equation to solve for x.
0.04x+0.05(15000-x) = 690 Perfom the indicated multiplication on the left side.
0.04x+750-0.05x = 690 Combine the x-terms on the left.
-0.01x+750 = 690 Subtract 750 from both sides of the equation.
-0.01x = - 60 Finally, divide both sides by -0.01
x = 6000
So Hakim invested $6,000 in the Term Deposit at 4% per annum.
He invested the remainder ($15,000-$6,000 = $9,000) in the Treasury Bill at 5% per annum.